# Omni-Mol: Exploring Universal Convergent Space for Omni-Molecular Tasks

### v2

@@ Abstract
Building generalist models has recently demonstrated remarkable capabilities in diverse scientific domains. Within the realm of molecular learning, several studies have explored unifying diverse tasks across diverse domains. However, negative conflicts and interference between molecules and knowledge from different domain may have a worse impact in threefold. First, conflicting molecular representations can lead to optimization difficulties for the models. Second, mixing and scaling up training data across diverse tasks is inherently challenging. Third, the computational cost of refined pretraining is prohibitively high. To address these limitations, this paper presents Omni-Mol, a scalable and unified LLM-based framework for direct instruction tuning. Omni-Mol builds on three key components to tackles conflicts: (1) a unified encoding mechanism for any task input; (2) an active-learning-driven data selection strategy that significantly reduces dataset size; (3) a novel design of the adaptive gradient stabilization module and anchor-and-reconcile MoE framework that ensures stable convergence. Experimentally, Omni-Mol achieves state-of-the-art performance across most of the 15 molecular tasks, demonstrates the presence of scaling laws in the molecular domain, and is supported by extensive ablation studies and analyses validating the effectiveness of its design. The dataset, code and weights of the powerful AI-driven chemistry generalist are open-sourced.
Machine Learning, ICML
[Uncaptioned image]Code:Omni-Mol-Code      [Uncaptioned image]Data:Omni-Mol-Data
@@ 1 Introduction
Large language models (LLMs), especially multimodal LLMs, have achieved significant breakthroughs in various scientific tasks due to their powerful representational capabilities and general reasoning abilities, spanning domains such as medicine (Jee et al., 2024; Zhou et al., 2024), chemistry (Boiko et al., 2023), and biology (Zhang et al., 2024b). This cutting-edge technology has also sparked an increasing number of studies exploring how to align molecular representation spaces with textual representation spaces (Cao et al., 2025; Chen et al., 2024; Fang et al., 2024; Cao et al., 2024). These works hold great promise to build powerful AI chemist for advancing molecule captioning, property/structure prediction, and text-conditioned de novo drug design.
@@
The first step in creating an AI chemist is to develop a generalist model with universal capabilities, enabling it to understand diverse molecular structures and their interactions under multiple chemical domains. Pioneering works, such as Text$+$Chem T5 (Christofidellis et al., 2023), introduces the first multi-domain, multi-task language model capable of unifying molecular and textual representations. Following this, PRESTO (Cao et al., 2024) further enhances performance by progressively improving multimodal LLMs through cross-modal alignment and multi-graph understanding. Similarly, in the field of general LLMs, the platonic representation (Huh et al., 2024) introduces the concept of the multitask scaling hypothesis, which suggests that as models are trained on an increasing number of tasks, they are driven to develop representations capable of addressing all tasks effectively. All of them highlight the potential of constructing the generalist model capable of handling a wide range of molecular tasks.
Figure 1: (Left) Illustration of the universal convergent representation space for omni-molecule tasks. (Right) Conflict between tasks from different domains makes vanilla models fail to converge.
However, we have yet to observe a model that achieves outstanding performances across as many tasks as possible, nor have we seen a clear trend toward scalability in this direction. For instance, InstructMol (Cao et al., 2025) attempts to scale up large language models but yields negligible gains, while PRESTO relies on a complex training strategy and requires extensive computational resources for pre-training. We propose that the fundamental challenge is conflict collapse, illustrated in Figure 1, which limits the emergence of truly generalist model in three key ways. First, potential conflicts may arise among various functional groups within a molecule and across the entire molecular structure, making it difficult to optimize the semantic relationships among different molecular representations. Second, data with conflicts from different domains often exhibit divergent distributions and interfere with each other, rendering it elusive to determine an ideal training data mixture. Third, the complexity of multi-task conflicts grows explosively as the volume of molecular data increases, requiring models with limited capacity to consume significantly greater resources in order to resolve these conflicts.
The recognition of the existing limitations naturally raises a pivotal question towards chemistry generalist model:
Is it possible to develop a generalist model that converges to a universal representation space for omni-molecular tasks?
This question drives us to develop Omni-Mol, a scalable and unified LLM-based framework for direct instruction tuning. Omni-Mol is the first generalist to effectively mitigate conflict collapse in three respects: (1) Omni-Mol proposes a unified encoding mechanism applicable to any task input, leveraging the most comprehensive instruction-following omni-molecular dataset to date, which comprises 1.8M samples across 15 tasks. (2) From the perspective of data, Omni-Mol employs an active learning-based dynamic data selection after recognizing that not all tasks are equally important. We significantly reduced the dataset size to 40% of its original volume while maintaining comparable performance with the full dataset. (3) From the perspective of architecture, our novel designed adaptive gradient stabilization successfully mitigate the rapid gradients growth caused by task conflicts (shown in Figure 2). Besides, our anchor-and-reconcile mixture-of-expert (MoE) architecture reduces interference by dynamically routing tasks to reconcile experts and capturing common knowledge through anchor experts.
@@
Extensive experiments show that Omni-Mol achieves significant improvements across most of the tasks, setting new state-of-the-art results among LLM-based models. Additionally, we observe that Omni-Mol scales effectively with increases in data volume, model size, and the number of tasks, indicating the model’s tremendous potential under larger computational budgets. Furthermore, by analyzing the representations of models trained with progressively more tasks, we discover that the representations become increasingly similar as the number of tasks grows. This provides robust evidence supporting our hypothesis of convergent space toward a universal molecular representation.
Figure 2: (Left) Gradient norm of unified training on 15 tasks. Due to the conflicts of multiple tasks, the gradient norm of InstructMol competes and shows a significant increase, while the gradient norm of Omni-Mol remains relatively stable. (Right) The scaling trend with task numbers on reagent prediction. As the number of tasks increases, Omni-Mol is benefited and consistently achieves better performances averagely, while InstructMol fails to scale up.
@@ 2.1 Molecular Foundation Models
Researchers are trying to leverage the world knowledge embedded in LLMs to build higher-quality molecular representations by fine-tuning on task-specific instructions. Mol-Instruction (Fang et al., 2024) pioneers the instruction fine-tuning dataset, demonstrating the potential of LLMs in molecular modeling. Subsequently, InstructMol (Cao et al., 2025) introduces 2D graph features of molecules based on SMILES (Weininger, 1988), showing that LLMs can also enhance performance by aligning and fine-tuning their understanding of graph-based features. Soon after, 3D-MoLM (Li et al., 2024b) explores the advantages of 3D molecular representations in multimodal LLMs, while HIGHT (Chen et al., 2024) investigates the impact of multi-level 2D graph features on molecular understanding. More recently, PRESTO (Cao et al., 2024) enhances LLMs’ comprehension of molecular-related knowledge through extensive domain-specific pretraining across eight tasks.
Figure 3: Overview of our proposed Omni-Mol, a scalable and unified LLM-based framework for direct instruction tuning.
@@ 2.2 Unified Generative Modeling
The GPT models (Brown et al., 2020; Achiam et al., 2023) have achieved unification across all text-based tasks through large-scale pretraining and instruction tuning. Subsequently, the community has successfully constructed models that can understand data from multiple modalities and simultaneously perform tasks related to different modalities by converting features from each modality into tokens (Alayrac et al., 2022; Li et al., 2022, 2023; Dai et al., 2023; Liu et al., 2024b). More recently, the community has also been exploring unified understanding and generation, allowing models not only to understand multimodal data but also to generate multimodal data (Zhu et al., 2023; Zheng et al., 2023a; Koh et al., 2024). This development is driving models towards convergence into a truly general-purpose model capable of solving all tasks. Huh et al. (2024) suggests that as models grow more powerful and general, their representations tend to converge, approaching a universal space that reflects the fundamental laws of the world. This insight inspires us to explore whether a universal convergent space also exists in the molecular domain.
@@ 3 Method
3.1 Overview
Omni-Mol is a multimodal LLM framework to handle $K$ diverse molecular tasks simultaneously. It comprises a language model, a graph encoder $f_{\mathcal{G}}$, and a projector $f_{p}$. The inputs include a text instruction $\mathbf{X}_{I}$, a SELFIES string $\mathbf{X}_{S}$, and the graph data $\mathbf{X}_{G}$ corresponding to the input molecules, where $\mathbf{X}_{G}$ is converted from $\mathbf{X}_{S}$ using RDKit. We model the response $\mathbf{Y}$ as the probability of the next token as:
$P(\mathbf{Y}|\mathbf{X}_{I},\mathbf{X}_{S},\mathbf{H}_{G})$ (1)
$=\prod_{i}P_{\theta}(\mathbf{Y}_{i}|\mathbf{X}_{I},\mathbf{X}_{S},\mathbf{H}_{G},\mathbf{Y}_{<i})$
where $\mathbf{H}_{G}=f_{p}(f_{\mathcal{G}}(\mathbf{X}_{G}))$, and $\theta$ is the parameter of the LLM. The graph encoder encodes the molecule graph into its representation $\mathbf{h_{g}}\in\mathbb{R}^{n\times d_{1}}$, where $n$ is the length of the representation, the projector then projects its dimension to the LLM’s hidden size and obtain $\mathbf{H}_{G}\in\mathbb{R}^{n\times d_{2}}$.
3.2 Unified Encoding for Instruction Tuning
Unified Input format. We collect and format the data of any task into the following structure.
  • Instruction $\mathbf{X}_{I}$: A brief, clear, and distinguishable instruction that tells the model what task to perform.
  • Input $\mathbf{X}_{S}$: A sequence of molecules represented using SELFIES, with different molecules separated by a dot.
  • Output $\mathbf{Y}$: The output corresponding to the task, which may be a number, a SELFIES representation, or a textual description, all in text string.
The data above will be processed according to a specific template, detailed information can be found in Appendix J.
Unified Encoding. To unify tokens from diverse tasks and modalities, and to enable parallel training on samples of varying lengths, we apply uniform padding to the mini-batch samples. First, padding tokens are added to the right side of the text input, and the batch is then passed through an embedding layer to obtain text embeddings. Next, since the number of atoms in each molecule varies, the number of graph tokens per sample differs as well. After inserting the graph tokens, we continue to add padding token embeddings to the right end of the sequence, ensuring the mini-batch becomes a well-formed tensor. We then generate an attention mask based on the padded tensor and assign an ‘ignore_index’ in the labels to prevent Omni-Mol from learning to generate padding tokens.
3.3 Active Learning-based Data Selection
To facilitate effective unified tuning, we systematically screen data from multiple tasks to minimize redundancy and conflicts, while determining an optimal mixing ratio. Actually, not all task-specific data is equally crucial. Inspired by Yu et al. (2024b), we employ an iterative task-centric data filtering approach to actively screen multiple datasets, substantially reducing training costs.
Given $K$ task collections $\mathcal{T}=\bigcup_{q=1}^{K}\mathcal{T}_{q}$, along with a total budget $\mathcal{B}$ and a maximum iteration count $J$. For the $j$-th iteration $\bigl{(}1\leq j\leq J\bigr{)}$, we select a data portion of size $\alpha_{j}=\mathcal{B}/J$. Assume the initial parameters of the model is $\theta_{0}$. For the $K$ subsets, we first initialize the distribution $\boldsymbol{\pi}^{0}=\bigl{(}\pi_{1}^{0},\,\pi_{2}^{0},\,\ldots,\,\pi_{K}^{0}\bigr{)},$ such that $\pi_{q}^{0}=1/K$, Then, we sample each task dataset $\mathcal{T}_{q}$ to a subset $\mathcal{T}_{q}^{(1)}$ according to the distribution and create the sampled training set $\mathcal{T}^{(1)}$. We then finetune $\theta_{0}$ on $\mathcal{T}^{(1)}$ for $M$ epochs, yielding the updated parameter $\theta_{1}$.
After obtaining $\theta_{1}$, we perform inference on the sampled dataset $\mathcal{T}^{(1)}$, we collect sample pair $(u_{p,+},u_{p,-})$ indexed by $p$, where $u_{p,+}$ is the ground truth and $u_{p,-}$ is the answer generated by the model. For each sample pair, we calculate the score of the sample by:
$\nu(u_{p,+},u_{p,-})=\max\{\rho(\cdot,\cdot)\}-\rho(u_{p,+}.u_{p,-})$ (2)
where $\rho(\cdot,\cdot)$ is a normalized metric function for each respective task. For each sample pair $(u_{p,+},u_{p,-})$ in the task dataset $\mathcal{T}_{q}$, we define the average rating as:
$\mu_{q}$ $=\frac{1}{\bigl{|}\mathcal{T}_{q}\bigr{|}}\sum_{(u_{p,+},u_{p,-})\in\mathcal{T}_{q}}\nu\bigl{(}u_{p,+},u_{p,-}\bigr{)}.$ (3)
A larger $\mu_{q}$ indicates higher complexity, implying that samples in this task are more instrumental for the model. At the end of the $j$-th iteration, we reweight the distribution via:
$\pi_{q}^{\prime(j)}=\frac{\mu_{q}}{\sum_{r=1}^{K}\mu_{r}}\pi_{q}^{(j-1)},\quad\pi_{q}^{(j)}=\frac{\pi_{q}^{\prime(j)}}{\sum_{r=1}^{K}\pi_{r}^{\prime(j)}}$ (4)
We then select an additional batch of size $\alpha_{j+1}=\mathcal{B}/J$ denoted $\mathcal{T}^{(j+1)}$, from the unused portion of the dataset. We concatenate it with the previously chosen data for further training. This sample–train–evaluate–reweight routine continues until $J$ total iterations are reached.
3.4 Stable & MoE Expanded Framework
Adaptive Gradient Stabilization. During unified training, each mini-batch consists of samples from several tasks. The resulting loss and gradients will be a combination of contributions from each task. Let $B_{k}$ denotes the number of samples for task $k$, and $T$ the sequence length. The total loss for the model can then be formulated as:
$\mathcal{L_{\text{total}}}=\frac{1}{B}\sum_{k=1}^{K}\sum_{b=1}^{B_{k}}\sum_{t=1}^{T_{k}}\mathcal{L}(m_{b,t}^{(k)},o_{b,t}^{(k)})$ (5)
where $m_{b,t}^{(k)}$ and $o_{b,t}^{(k)}$ are the predicted logits and label for batch element $b$ at time step $t$ in task $k$.
Actually, we observe a noticeable increase in gradient norms (shown in Figure 2), which leads to training instability and hindered faster loss convergence. We attribute the main cause of these divergences to issues arising from the softmax operation when handling tasks drawn from different domains with substantially varied entropy, stemming from softmax’s translation invariance (i.e., $softmax\left(z\right)=softmax\left(z+c\right)$). Since all model parameters are shared among multiple tasks, each task competes by incrementally growing its norms. while this is not immediately detrimental, it leads to divergence once norms extend beyond the effective range of bf16. To mitigate this problem, we employ an adaptive coefficient $\gamma_{\theta}=\frac{\alpha_{\theta}}{||r||_{p}}+\beta_{\theta}$ during the parameter-efficient fine-tuning with LoRA (Hu et al., 2021), where $\alpha_{\theta}$ and $\beta_{\theta}$ are learnable variables as scaling factors and $r$ represents the rank of LoRA. Through this mechanism, gradients can be adaptively stabilized on data $\mathcal{D}$:
$\nabla_{\boldsymbol{\Delta{\mathcal{W}}}}=\frac{\partial\mathcal{L}_{total}\left(\mathcal{D};\boldsymbol{\mathcal{W}}_{0}+\gamma_{\theta}\cdot\boldsymbol{\Delta{\mathcal{W}}}\right)}{\partial\boldsymbol{\Delta{\mathcal{W}}}}$ (6)
where $\Delta\mathcal{W}$ is the updated parameters to the pre-trained $\mathcal{W}_{0}$. More details can be seen in Appendix C.3.
Anchor-and-Reconcile Experts Expansion. Omni-Mol needs to learn a wide range of different tasks and handle multiple modalities, including graph features, text, and SELFIES. While SELFIES is treated as regular text input to the LLM, it inherently differs significantly from natural language semantics, requiring the model to separately learn how to understand and generate SELFIES expressions.
We aim for the model to simultaneously learn general knowledge while also differentiating for different modalities and tasks. Hence, we borrow the idea of MoE (Dai et al., 2024) and perform upcycling (Komatsuzaki et al., 2023; Lin et al., 2024). We first construct $\mathcal{N}$ reconcile experts, each targeting specialized knowledge areas, and dynamically balances conflicting signals among these experts to effectively mitigate task-level conflicts. Besides, we introduce an additional anchor expert to learn the common knowledge that underpins fundamental understanding across tasks, by consistently capturing and aligning shared features to maintain a stable global representation.
To be specific, for a regular decoder layer $l=1\dots L$ from a pre-trained LLM,
$h^{\prime}_{l}=h_{l-1}+\text{MHA}_{\phi}(\text{LN}(h_{l-1}))$ (7)
$h_{l}=h^{\prime}_{l}+\text{FFN}_{\gamma}(\text{LN}(h^{\prime}_{l}))$
where $\phi$ and $\gamma$ are parameters of the pre-trained LLM. We convert the decoder layer into:

### v3

# Omni-Mol: Multitask Molecular Model for Any-to-any Modalities

@@ Abstract
In the molecular domain, numerous studies have explored the use of multimodal large language models (LLMs) to construct a general-purpose, multi-task molecular model. However, these efforts are still far from achieving a truly universal molecular model. We identify three key challenges in this endeavor: (1) Existing molecular task datasets are typically small in scale and lack comprehensive domain coverage. (2) Tasks from different molecular subfields are difficult to effectively learn jointly through LLMs due to significant distributional shifts and competition among tasks, which introduces instability in the learning process. (3) Both inter-task and intra-task molecular representations demand different intrinsic dimensions in the language space, making it challenging to balance between redundancy and insufficiency in language model representations. To address these challenges, we innovatively categorize existing small-molecule tasks into four types: Mol2Mol, Mol2Text, Mol2Num, and Text2Mol. We then collect a dataset encompassing over 16 tasks with more than 1.4 million samples, making it the largest molecular instruction-tuning dataset to date. Leveraging the extensive pretraining of LLMs on existing chemical literature, we propose a novel multimodal LLM framework, named Omni-Mol, which unifies all small-molecule tasks and supports both molecular generation and understanding. The core of Omni-Mol is our proposed MoGE, which dynamically adapts to the intrinsic rank of different tasks. This mixture-of-experts architecture enhances the model’s ability to handle diverse tasks and modalities effectively. Our model achieves unified instruction tuning across 16 tasks and attains state-of-the-art performance on 13 of them. Extensive experiments further demonstrate the scalability and versatility of Omni-Mol.
[Uncaptioned image]GitHub:Omni-Mol-Code     [Uncaptioned image]HuggingFace:Omni-Mol-Data&Weight
@@ 1 Introduction
Large language models (LLMs), especially multimodal LLMs, have achieved significant breakthroughs in various scientific tasks due to their powerful representational capabilities and general reasoning abilities, spanning domains such as medicine jee2024automated ; zhou2024pre , chemistry boiko2023autonomous , and biology zhang2024multimodal . This cutting-edge technology has also sparked an increasing number of studies exploring how to align molecular representation spaces with textual representation spaces cao2023instructmol ; chen2024hight ; fang2024molinstructions ; cao-etal-2024-presto ; hu2024exploring . These works hold great promise to build powerful AI chemists for advancing molecule captioning, property/structure prediction, and text-conditioned de novo drug design.
@@
The first step in creating an AI chemist is to develop a generalist model with universal capabilities, enabling it to understand diverse molecular structures and their interactions under multiple chemical domains. Pioneering works, such as Text++Chem T5 christofidellis2023unifying , introduce the first multi-domain, multi-task language model capable of unifying molecular and textual representations. Following this, the recent state-of-the-art PRESTO cao-etal-2024-presto further enhances performance by progressively improving multimodal LLMs through cross-modal alignment and multi-graph understanding.
From this trend, it can be seen that the community is pursuing a "one-model-fits-all" zhu2023minigpt4 ; zheng2023minigpt5 paradigm, rather than using models like InstructMol cao2023instructmol that rely on different Low-Rank Adapters(LoRA) hu2021lora to learn different tasks.
Benchmark
1. No LLM Pre-training
2. One model fits all
3. 3D adaptive ability
4. Data scaling
5. Parameter scaling
6. Task type
InstructMol cao2023instructmol 3
HIGHT chen2024hight 3
PRESTO cao-etal-2024-presto 2
3D-MoLM li2024towards 2
ReactXT liu2024reactxt 2
Omni-Mol 4
Table 1: A comprehensive comparison between Omni-Mol and other molecular LLMs.
However, existing approaches remain far from achieving a truly general-purpose molecular model. For instance, while PRESTO can generate molecules given molecular inputs or predict properties based on molecular structures, it does not support tasks such as describing molecules or designing molecules according to specified textual requirements. To date, it is rare to find a model that supports a sufficiently broad range of task types under a unified “one-model-fits-all” framework. We identify three main challenges in constructing such a universal molecular model. First, existing molecular task datasets are generally small in scale and lack coverage across diverse domains. Second, molecular tasks from different domains exhibit significant distributional discrepancies, making it difficult for LLMs to learn effectively and stably across tasks. Finally, both intra-task molecular instances and inter-task representations differ in their intrinsic dimensionality within language space, making it challenging for the model to balance redundancy and insufficiency. These issues impede the development of a general-purpose AI expert for molecular tasks.
In this paper, we seek the answer to the following question:
Is it possible to develop a generalist molecular LLM capable of effectively learning across diverse task domains?
This question drives us to develop Omni-Mol, a scalable and general-purpose Multimodal LLM-based framework for unified molecular understanding and generation. Omni-Mol provides four key innovations. (1) we conduct a comprehensive investigation of small molecule tasks and innovatively categorize these tasks according to their input-output modalities into four types: Mol2Mol, Mol2Text, Mol2Num, and Text2Mol. Subsequently, we construct the Omni-Mol dataset, which comprises over 1.4 million samples and represents the most extensive instruction-tuning dataset for small molecule tasks to date. (2) Leveraging this dataset, we propose a unified instruction tuning paradigm and build the most comprehensive general-purpose multimodal molecular LLM based on LLaMA 3 dubey2024llama for the first time. (3) To address the challenge of varying intrinsic dimensions across different domains and tasks, we propose Gradient Adaptive LoRA (GAL), a novel adaptive mechanism that extends existing LoRAs hu2021lora ; ding2023parameter ; wang2023parameter ; zhai2023parameter ; yu2024visual ; wang2024lion to better handle multi-task learning scenarios. GAL mitigates conflicts that arise when standard LoRA struggles to accommodate dynamically shifting intrinsic dimensions during training. (4) To further manage inter-task and cross-modal interference, we adopt a Mixture-of-Experts (MoE) framework to develop a Mixture-of-GAL-Experts (MoGE) fine-tuning strategy. By integrating shared experts and routed experts, our model is capable of both robustly capturing general knowledge and differentiating across diverse tasks.
@@ 6. Task type
Comprehensive experiments on our datasets show that Omni-Mol achieves significant improvements across 13 tasks simultaneously, setting new state-of-the-art results among both finetuned open-source LLMs and in-context learned closed-source LLMs. Additionally, we observe that Omni-Mol scales effectively with increases in data volume and model size, indicating the model’s tremendous potential under larger computational budgets. Furthermore, by analyzing the representations of models trained on progressively more tasks, we discover that the representations become increasingly similar as the number of tasks grows. This provides robust evidence that the model is learning general representations effectively. We hope our dataset and model can pave the way for the community to build more powerful generalist AI chemists.
@@ 2.1 Molecular Foundation Models
Researchers are trying to leverage the world knowledge embedded in LLMs to build higher-quality molecular representations by fine-tuning on task-specific instructions. Mol-Instruction fang2024molinstructions pioneers the instruction fine-tuning dataset, demonstrating the potential of LLMs in molecular modeling. Subsequently, InstructMol cao2023instructmol introduces 2D graph features of molecules based on SMILES weininger1988smiles , showing that LLMs can also enhance performance by aligning and fine-tuning their understanding of graph-based features. Soon after, 3D-MoLM li2024towards explores the advantages of 3D molecular representations in multimodal LLMs, while HIGHT chen2024hight investigates the impact of multi-level 2D graph features on molecular understanding. More recently, PRESTO cao-etal-2024-presto enhances LLMs’ comprehension of molecular-related knowledge through extensive domain-specific pretraining across eight tasks.
@@ 2.2 Unified Generative Modeling
The GPT models brown2020language ; achiam2023gpt have achieved unification across all text-based tasks through large-scale pretraining and instruction tuning. Subsequently, the community has successfully constructed models that can understand data from multiple modalities and simultaneously perform tasks related to different modalities by converting features from each modality into tokens alayrac2022flamingo ; li2022blip ; li2023blip ; dai2023instructblip ; liu2024visual . More recently, the community has also been exploring unified understanding and generation, allowing models not only to understand multimodal data but also to generate multimodal data zhu2023minigpt4 ; zheng2023minigpt5 ; koh2024generating . This development is driving models towards convergence into a truly general-purpose model capable of solving all tasks. huh2024platonic suggests that as models grow more powerful and general, their representations tend to converge, approaching a universal space that reflects the fundamental laws of the world. This insight inspires us to explore whether a multi-task generalist also exists in the molecular domain.
Task #Samples
Mol2Mol 688,725
Forward 125,384
Reagent 125,384
Retrosyn 129,684
Solvent 74,892
Catalyst 11,094
MolEdit 222,287
Task #Samples
Mol2Num 412,231
HO.LU. 263,100
Weight 13,979
TPSA 13,979
LogP 12,458
Yield 9,715
Task #Samples
Mol2Text 247,652
Experiment 147,788
DQA 66,885
Molcap 32,979
Task #Samples
Text2Mol 72,559
I2S 57,575
MolDesign 14,984
Table 2: Task statistics, HO.LU.: HOMO-LUMO prediction, I2S: IUPAC2SELFIES. The total molecule data volume reaches 1.42 million.
(a) Domain and tasks
(b) Molecule statistics
Figure 1: The sunburst chart of Omni-Mol dataset tasks and the statistics of the molecules. Our dataset exhibits substantial diversity across several molecular attributes, including atom count, ring count, molecular weight, and Bertz complexity.
3 Overview
Omni-Mol is a multimodal LLM framework to handle $K$ understanding and generative molecular tasks simultaneously. It comprises a language model, a graph encoder $f_{\mathcal{G}}$, and a projector $f_{p}$. The inputs include a text instruction $\mathbf{X}_{I}$, a SELFIES string $\mathbf{X}_{S}$, and the graph data $\mathbf{X}_{G}$ corresponding to the input molecules, where $\mathbf{X}_{G}$ is converted from $\mathbf{X}_{S}$ using RDKit. We model the response $\mathbf{Y}$ as the probability of the next token as:
$P(\mathbf{Y}|\mathbf{X}_{I},\mathbf{X}_{S},\mathbf{H}_{G})=\prod_{i}P_{\theta}(\mathbf{Y}_{i}|\mathbf{X}_{I},\mathbf{X}_{S},\mathbf{H}_{G},\mathbf{Y}_{<i})$ (1)
where $\mathbf{H}_{G}=f_{p}(f_{\mathcal{G}}(\mathbf{X}_{G}))$, and $\theta$ is the parameter of the LLM. The graph encoder encodes the molecule graph into its representation $\mathbf{h_{g}}\in\mathbb{R}^{n\times d_{1}}$, where $n$ is the length of the representation, the projector then projects its dimension to the LLM’s hidden size and obtain $\mathbf{H}_{G}\in\mathbb{R}^{n\times d_{2}}$. The overview is shown in Figure 2. The complete multimodal architecture details are in Appendix C.
4 Omni-Mol Data Curation
The first step in constructing Omni-Mol is the collection of comprehensive and diverse data. In our review of existing work, we identify a wide range of chemical tasks, whose inputs and outputs can generally be categorized into the following modalities: Text, 1D molecular sequences, and Tabular numerical value. For instance, a typical chemical reaction task involves mapping input molecules to another molecule, which falls under the category of molecule-to-molecule tasks. Some studies cao2023instructmol ; cao-etal-2024-presto ; chen2024hight incorporate graph neural networks (GNNs) to encode molecular graph information as input; however, graph features are typically not used as outputs.
After a systematic review, we categorize the tasks into the following four major types: (1) Mol2Mol, (2) Mol2Num, (3) Mol2Text, and (4) Text2Mol. We observe that existing works are not yet capable of achieving learning across arbitrary modality pairs. For example, PRESTO covers Mol2Num and Mol2Mol tasks but lacks support for Mol2Text and Text2Mol. Similarly, InstructMol and 3D-MoLM include Mol2Num, Mol2Mol, and Mol2Text tasks, but do not support Text2Mol. Omni-Mol will be trained across all four types of tasks.
Data format. To this end, we construct a unified instruction-tuning dataset, standardizing the data format as follows: (1) A unique, clear, and concise instruction; (2) A 1D representation of the molecule (no molecule input for Text2Mol tasks); (3) The corresponding task output.
SELFIES v.s. SMILES. Both SELFIES krenn2020self and SMILES weininger1988smiles are 1D modalities for representing molecules as text. In the Omni-Mol Dataset, we opted for the SELFIES representation. We observed that SMILES strings may fail to be recognized by RDKit landrum2013rdkit ; this issue is particularly pronounced in SMILES generated by LLMs. In contrast, a syntactically correct SELFIES string can be robustly decoded into a valid molecule krenn2020self . This reliable decoding is necessary to guarantee the conversion to a 2D graph, which in turn ensures the generation of a 3D graph. These 2D and 3D graph representations are critical for downstream tasks such as molecular docking and conformer generation.
Preprocessing. After collecting the data, we perform a series of preprocessing steps and establish comprehensive metrics for molecular LLMs on the selected task. We also look into tasks that are similar to each other and remove samples that are potentially data leakage. To further investigate the understanding of 3D molecules by LLMs, following 3D-MOIT li2024towards , we preprocess the Omni-Mol data using RDKit to obtain the 3D representation graphs of the molecules. Detailed information can be found in Appendix B.
Unified Encoding. We encode data from different modalities uniformly into tokens. For molecular 1D representations, textual data, and tabular numerical data, we convert them into character sequences and tokenize them using the tokenizer and word embeddings of the LLM. For molecular graph representations, we utilize the node embeddings generated by a GNN as input tokens. Since the lengths of molecular sequences and graph embeddings vary, we apply padding to standardize their shapes into a uniform tensor format, thereby enabling parallel training.
Dataset Statistic. We conducted a comprehensive statistical analysis of the entire dataset, including metrics such as atom count and ring count. The task types along with the names of individual tasks in Figure 1. The final Omni-Mol dataset contains over 1.4 million samples, providing ample training fuel for the development of Omni-Mol. The detailed information on each task and the chat template can be found in Appendix B. Moreover, we introduce MolEdit as a novel task within our collected data. As the original data was not formatted for instruction-following, we designed and authored specific instructions for molecular editing.
@@ 5 Method
Figure 2: Overview of our proposed Omni-Mol, a scalable and general-purpose molecular LLM-based framework for both understanding and generation.
5.1 Gradient Adaptive LoRA